1. How efficient is a machine that takes in 24000J of energy and produces 8000J of useful energy?


  1. A machine which is 70% efficient has a power input of 360W.  What is the power output?


  1. A motor which is 60% efficient has an output power of 480 W. What is its input power?


  1. An electric motor rated at 500 W runs for 2 minutes and does 45 kJ of work.


a)     What is its input energy?

b)     What is its efficiency?


  1. What is the efficiency of a motor rated at 750 W which can lift a 25 kg load to a height of 4 m in 2 s?
1
Expert's answer
2021-05-12T08:38:24-0400

1.

"\eta=\frac{Q_o}{Q_i}=0.33=33\%."

2.

"P_o=\eta P_i=252~W."

3.

"P_i=\frac{P_o}{\eta}=800~W."

4a.

"A_i=Pt=60~kJ."

4b.

"\eta=\frac{A_o}{Pt}=0.75=75\%."

5.

"\eta=\frac{A}{Pt}=\frac{mgh}{Pt}=0.67=67\%."